OpenStudy (anonymous):

f(x)=x^(3)+3x, find each value a. f(2+h)-f(2) b. f(1/4) c. f(√2) d f(1+h) e. f(1) f. f(1+h)-f(1)

OpenStudy (anonymous):

f(2+h) = substitute (2+h) in place of x in the function f(x) and then solve

jhonyy9 (jhonyy9):

you need make calculi how was wrote above but looke the f(2)=2^3 +3*2 = 8 +6 = 14 so you need calculi f(2+h) = the same how i have calculed above f(2) ok ?

OpenStudy (anonymous):

thats to say f(2+h)=(2+h)^3+3*2+h right?

OpenStudy (anonymous):

h^3+6h^2+12h+8 - 14

jhonyy9 (jhonyy9):

so above you need multiplie 3 with (2+h) and not just with 2 ok ?

OpenStudy (anonymous):

h^3+6h^2+12h+8+6+3h -14 =h^3+6h^2+15h right?

jhonyy9 (jhonyy9):

i think yes

OpenStudy (anonymous):

tnx and what about f(1/4)?

jhonyy9 (jhonyy9):

write it on the open question columne please